$10,000.00 dollars compounded daily at 1.5% interest is how much after 12 months?
Answer by Theo(13342)

$10,000 compounded daily for 12 months is equal to:
10,000 * (1+x)^365,...
if we assume there are 365 days in a year.
this does not take into account leap years where there are 366 days in a year.
your equation is:
f = p * (1+x)^n
f = future value
p = present value
x = interest rate per time period.
n = number of time periods.
the equation becomes:
f = 10,000 * (1+x)^365
we know that x = 1.5%
the decimal equivalent of 1.5% is .015
that, presumably, is the interest rate per time period which means that it is the interest rate per day.
our equation becomes:
f = 10,000 * (1.015)^365
solve for f using your calculator to get:
f = 2,291,423.807
if the interest rate is the annual interest rate rather than the daily interest rate, then the formula is changed as follows:
interest rate per year = .015
interest rate per day = .015/365 = .000041096 (rounded to 9 decimal places)
your formula becomes:
f = 10,000 * (1.000041096)^365
solve for f using your calculator to get:
f = 10,151.12792
big difference.